If 26 V is applied across the whole network of Fig. 19-52,calculate the voltage across each capacitor.
Answer #1
Given
   capcitance C1 = 3.0 μF=3.0*10-6F
   capacitance C2 =  4.0 μF=4.0*10-6F
   capcitance C3 = 2 μF =2*10-6F
   voltage V = 26V
------------------------------------------------------------------------------
   
   capacitance C1 and C2 are connected isseries
the equavalent capacitor Ceq = C1C2 / ( C1 +C2)
                                          = 1.71*10-6F
   Hence charge on the capacitor
                  Q = Ceq V
                    = 44.57*10-6C   
    voltage across the capacitorC1  
   
                   V1 = Q /C1
    voltage across the capacitorC2  
   
                   V2 =V- V1
   substitute the values for V1 and V2
   capacitor C3 is parallel to C1 and C2 so thatvoltage across C3 is 26V
Answer #2
capacitors C1 and C2 are connected in series:


1/Ceq=1/c_1 +1/c_2
1/Ceq=1/3+1/4
1/Ceq=.5833
Ceq=1.71